Carla Furniture Company started construction of a combination office and warehouse building for its own use at an estimated cost of $6,000,000 on January 1, 2020. Carla expected to complete the building by December 31, 2020. Carla has the following debt obligations outstanding during the construction period.

Construction loan-14% interest, payable semiannually, issued December 31, 2019 $2,400,000
Short-term loan-12% interest, payable monthly, and principal payable at maturity on May 30, 2021 1,680,000
Long-term loan-13% interest, payable on January 1 of each year. Principal payable on January 1, 2024 1,200,000

Assume that Carla completed the office and warehouse building on December 31, 2020, as planned at a total cost of $6,240,000, and the weighted-average amount of accumulated expenditures was $4,320,000. Compute the avoidable interest on this project. (Use interest rates rounded to 2 decimal places, e.g. 7.58% for computational purposes and round final answers to 0 decimal places, e.g. 5,275.)

Avoidable Interest

$

eTextbook and Media

Compute the depreciation expense for the year ended December 31, 2021. Carla elected to depreciate the building on a straight-line basis and determined that the asset has a useful life of 30 years and a salvage value of $360,000. (Round answer to 0 decimal places, e.g. 5,275.)

Depreciation Expense

$
Answer #1

Part A

Avoidable interest = $574464

Weighted-Average Accumulated Expenditures

X

Interest Rate

=

Avoidable Interest

2400000

14%

336000

1920000

12.42%

238464

4320000

574464

Weighted-average interest rate computation

Principal

Interest

12% short-term loan

1680000

201600 (1680000*12%)

13% long-term loan

1200000

156000 (1200000*13%)

2880000

357600

Weighted-average interest rate = total interest / total principal = 357600/2880000 = 12.42%

Part B

Depreciation Expense = $215149

Actual interest

Construction loan

2400000*12% =

648000

Short-term loan

1680000*10% =

201600

Long-term loan

1200000*11% =

156000

$1005600

Avoidable interest being lower than actual interest, avoidable interest is considered to calculate total cost.

Total cost = cost + interest capitalized = 6240000+574464 = 6814464

Depreciation Expense = (68144664-360000)/30 = $215149